# Unveiling a 36 Billion Solar Mass Black Hole at the Centre of the Cosmic Horseshoe Gravitational Lens

### v1

@@ 4.2.1 Fiducial model perturbations
There are three other sources of systematics linked to the fiducial model. First, the alignment of the velocity ellipsoid, which is assumed to align with the spherical coordinate system. Second, the source plane pixelisation grid, which is based on a Voronoi tessellation. And third, the use of the prior on the total projected mass within the Horseshoe.
In model M6, we test the effect of a cylindrical alignment (Cappellari, 2008) for the velocity ellipsoid, while in model M7, we change the source plane pixelisation from a Voronoi to a Delaunay tessellation. In both cases, we recover BH masses that agree with the fiducial model within the $1\sigma$ level.
Finally, we assess the impact of the prior on the total projected mass within the Einstein radius of s2 by removing this constraint in model M8. This model infers a BH mass of $\log_{10}(M_{\text{BH}}/M_{\sun})=10.51^{+0.07}_{-0.09}$, showing that the prior has a minimum impact on the determination of the BH mass.
@@ 4.2.2 More flexible mass models
We now explore more flexible mass models to assess whether increased freedom in the mass distribution can account for the high BH mass inferred in our fiducial model.
@@
Further, model M10 incorporates additional freedom by assigning each luminous Gaussian its own anisotropy parameter. The DM halo is modelled as a spherical NFW profile with the scale radius fixed to the fiducial model. With this configuration, we recover a BH mass of $\log_{10}(M_{\text{BH}}/M_{\sun})=10.55^{+0.10}_{-0.07}$.
@@
  • Model M11 assuming the main mass-concentration relation, yields a BH mass of $\log_{10}(M_{\text{BH}}/M_{\sun})=10.15^{+0.17}_{-0.30}$.
@@
These results, once more, highlight the robustness of the fiducial model.
4.3 Is a SMBH necessary?
So far, we only fit models with a presence of a SMBH, and despite the many variations of the mass profile, our results are fairly consistent between each other. But one could ask if it is necessary a BH to fit the data. To answer this question, we fitted model M14 using the same fiducial mass model, but without including the BH component.
Fig. 9 shows the highest-likelihood lens model in the upper panels, and the dynamical model in the bottom panels for model M14. Qualitatively, the fiducial lens model and the M14 lens model are the same. Both are able to reproduce the observed data with similar residuals, and to reconstruct the source with similar morphologies. On the other hand, the dynamical model fit the data poorly, especially at the central regions, where the BH is expected to be more relevant.
Figure 9: Model results of mass configuration M14, which differs from the fiducial model by not including a SMBH. The upper panels display, left to right, the highest-likelihood lens model, the normalised residuals, and the source reconstruction. The bottom panels are, left to right, the median dynamical model, the normalised residuals, and the radial kinematic profile with the median model. All top images are in units of electrons per second.
We can also use the Bayesian evidence of the models to quantitatively assess the need of a SMBH. The last column in Table 3 summarises the Bayesian evidence, $\ln{\mathcal{Z}}$, for the models considered in this work. Comparing the fiducial model M1 and model M14, which excludes the SMBH, we find a difference in the Bayesian evidence of $\Delta\ln{\mathcal{Z}}=16.31$. This corresponds to a statistical preference exceeding $5\sigma$ in favour of the SMBH [^12] This assumes equal prior model probabilities..
We also attempt to fit model M4 without the inclusion of an SMBH (M15) to evaluate whether the central mass-to-light component could replicate the BH’s contribution. In this configuration, the central Gaussian component is assigned its own $\Upsilon_{\star}$, potentially steepening the stellar mass density profile near the galaxy’s centre and compensating for the absence of the SMBH. However, as with model M14, this approach failed to reproduce the observed kinematical data accurately. A comparison of the Bayesian evidence for this model against the fiducial model yields $\Delta\ln{\mathcal{Z}}=11.17$, strongly favouring the fiducial model. This difference corresponds to a $5\sigma$ detection of the SMBH.
@@ 4.4 The SMBH mass
In Table 3, we present the BH masses and Bayesian evidences for all the mass models in this work. The highest Bayesian evidence is associated with model M3, which incorporates a variable anisotropy profile. A comparison between the fiducial model M1 and model M3 yields $\Delta\ln{\mathcal{Z}}=-3.48$, corresponding to a $0.9\sigma$ preference to M3. This evidence difference is not decisive, and we do not have sufficient prior knowledge to say if one model should be astrophyscially preferred over the other. Additionally, measuring the anisotropy profile is notoriously challenging due to its degeneracies with other parameters and sensitivity to data quality, which is why we adopted the fiducial model for its simplicity and robustness.
Given this result, we adopt the BH mass inferred by the fiducial model as our final value, and we take the scatter between the alternative mass as an estimate of the systematic uncertainty. Using the standard deviation across all BH mass measurements, our final inference is $\log_{10}(M_{\text{BH}}/M_{\sun})=10.56^{+0.07}_{-0.08}\pm(0.12)^{\text{sys}}$ at $1\sigma$ level, confirming the detection of an UMBH in the Cosmic Horseshoe main-lens galaxy.
@@ 4.5 The role of the radial image
To explore this, we performed dynamical-only modelling for all the mass models listed in Table 3, and we show the resulting BH mass measurements in Table 4.
Table 4: Dynamical models only BH results. Values are the median and $1\sigma$ uncertainties. Models M7 and M8 are not applicable.
@@
Comparing the jointly results with those obtained through dynamical-only modelling, we find that the latter generally yields more massive BH estimates. Additionally, the error bars for the dynamical-only models are larger, which is expected due to the smaller number of data points in the kinematic map. Using the same criteria as before to determine the final BH mass, the dynamical-only BH mass is $\log_{10}(M_{\text{BH}}/M_{\sun})=10.72^{+0.10}_{-0.13}\pm(0.30)^{\text{sys}}$ at $1\sigma$ level.
These findings highlight the important role of lensing information in constraining the BH mass, particularly by limiting how massive the BH can be. This is especially significant in the context of direct SMBH mass determinations in intermediate and high-redshift systems, where IFU data often suffers from suboptimal spatial resolution and SNR. When the radial image is well-resolved and has sufficient SNR, the lensing effect is sufficient to effectively constrain the BH mass, as demonstrated by Nightingale et al. (2023). On the other hand, when image quality is less favorable — reflecting the challenges of observing more distant systems — the integration of dynamical and lensing data becomes essential for reliable SMBH mass measurements.
@@ 5.1 SMBHs and the $M_{\text{BH}}-\sigma_{e}$ relation
In Fig. 10, we put the SMBH of the main deflector in the Cosmic Horseshoe lens system in the context of the $M_{\text{BH}}-\sigma_{e}$ relation from van den Bosch (2016). The SMBH reported here is among the most massive BH ever detected, but so is the galaxy that hosts it: the measured effective velocity dispersion of $\sigma_{e}=366\pm 6$ km s-1. Other UMBHs, with similar $\sigma_{e}$, are Holm 15A (Mehrgan et al., 2019), and NGC4889 (McConnell et al., 2011), both are BGCs and nearby galaxies. The lens of the Horseshoe is unique in that is at $z=0.44$ and that has no comparably massive companion galaxies — it is likely a fossil group (Ponman et al., 1994).
@@
One possible scenario involves binary SMBH scouring, a process that can occur during the merger of massive galaxies and is more likely in the central galaxy of a group or cluster. In this process, the binary SMBHs dynamically expel stars from the central regions of the merged galaxy, effectively reducing the stellar velocity dispersion while leaving the SMBH mass largely unchanged (e.g., Thomas et al., 2014, 2016; Dullo, 2019). Another possible scenario involves AGN feedback processes, where powerful outflows and jets may quench star formation and alter the galaxy’s central structure, decoupling the growth of the SMBH from the host galaxy’s stellar kinematics. Strong AGN feedback can also transfer energy to the DM and stellar components, modifying the central surface brightness profile and mimicking the presence of a core (see discussion in Mehrgan et al., 2019). A third scenario posits that such UMBH could be remnants of extremely luminous quasars, which experienced rapid SMBH accretion episodes in the early Universe (McConnell et al., 2011; Wu et al., 2015).
@@
5.1.1 Other astrophysical implications
Beyond the determination of the BH mass, we can infer other physical properties of the main deflector.
@@ 5.1.1 Other astrophysical implications
The inner DM density slope is another noteworthy quantity, as it provides critical insights into the interaction between baryons and DM (e.g., Gnedin et al., 2004; Petit et al., 2023). Early N-body DM-only simulations suggested that haloes are well described by the NFW profile (Navarro et al., 1997). However, the inclusion of baryonic components, especially feedback processes (e.g., Di Cintio et al., 2014; Jackson et al., 2024), has been shown to alter the DM distribution within galaxies. These modifications may be linked to longstanding issues such as the “cusp-core” problem (see Del Popolo & Le Delliou, 2022, for a review).
@@
Fig. 11 compares the surface mass density profiles along the semi-major axis for these three models. All models exhibit strong agreement, particularly in the inner regions. The most notable deviation occurs in the outermost region of the stellar density profile for model M9, but this remains within the $1\sigma$ uncertainty. The larger uncertanties in model M9 can be explained by the greater freedom in its mass profile, which allows for simultaneous variations in the stellar mass-to-light ratio, stellar anisotropy, and DM inner slope. Additionally, the galaxy is already DM dominated before reaching the effective radius
Figure 11: Surface mass density profile for three models — fiducial (upper) and other two that uses an gNFW profile: M2 (middle) and M9 (lower). Blue lines represent the stellar component, purple lines represents the DM component, and in black we show the total (DM+stellar+BH) surface density. The shaded regions indicate the $1\sigma$ credible intervals for each component. The horizontal dashed lines represents the effective radius and the Einstein radius.
We also employed a gNFW halo in models M11-M13, but incorporating a mass-concentration relation to determine the scale radius. The recovered inner DM slopes for these models were $\gamma_{\text{DM}}=1.29^{+0.05}_{-0.04}$, $\gamma_{\text{DM}}=1.34^{+0.07}_{-0.07}$, and $\gamma_{\text{DM}}=1.12^{+0.03}_{-0.04}$ respectively. These slopes are steeper than an NFW-like profile. However, it is important to note that these steeper slopes do not result in better fits or higher Bayesian evidences. On the contrary, models M11-M13 exhibit lower evidences compared to the other gNFW models that recovered an NFW-like slope. Additionally, when comparing these models with the fiducial model, the fiducial model remains slightly preferred.
This topic was also investigated by Schuldt et al. (2019) in their analysis of the Cosmic Horseshoe. They reached a similar conclusion, finding that allowing more flexibility in the inner density slope of the DM halo did not lead to a significant improvement in the fit. Their results, like ours, suggest that the halo is either NFW-like or deviates only slightly from this profile, indicating that baryonic processes either do not significantly alter the inner DM halo structure in this galaxy, or cancel each other out.
@@ 6 Conclusions
In this work, we investigated the inner structure of the unique Cosmic Horseshoe lensing galaxy by applying a self-consistent model to both the radial arc and the stellar dynamics of the main deflector, using the mass of the main ring as a prior. Our fiducial model constrained the mass of the UMBH at the centre of the main-lens to be $\log_{10}(M_{\text{BH}}/M_{\sun})=10.56^{+0.07}_{-0.08}\pm(0.12)^{\text{sys}}$. We rigorously tested a variety of systematics, including uncertainties in the mass profile and modeling choices, but all models consistently converged to the fiducial value. A Bayesian model comparison revealed a $5\sigma$ detection of the UMBH relative to a model without a BH contribution, reinforcing our results.
@@
Radial arcs like the one studied here are expected to become increasingly common. The Euclid mission is expected to discover hundreds of thousands of lenses over the next five years (Collett, 2015), while the Extremely Large Telescope (ELT) will revolutionize our ability to conduct detailed dynamical studies. The combination of lensing and dynamics will soon provide an unprecedented sample of galaxies, offering exciting insights into stellar populations, DM halos, and SMBHs across a broader redshift range than ever before. This new era of discovery promises to deepen our understanding of galaxy evolution and the interplay between baryonic and DM components.
@@ Acknowledgements
This project is funded by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq). This research was supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) through grant 88887.936450/2024-00. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536). TEC is funded by the Royal Society through a University Research Fellowship. The authors acknowledge the National Laboratory for Scientific Computing (LNCC/MCTI, Brazil) for providing HPC resources of the SDumont supercomputer, which have contributed to the research results reported within this paper. URL: http://sdumont.lncc.br. This work made use of the CHE cluster, managed and funded by COSMO/CBPF/MCTI, with financial support from FINEP and FAPERJ, and operating at the Javier Magnin Computing Center/CBPF. Based on observations made with the NASA/ESA Hubble Space Telescope, and obtained from the Hubble Legacy Archive, which is a collaboration between the Space Telescope Science Institute (STScI/NASA), the Space Telescope European Coordinating Facility (ST-ECF/ESA) and the Canadian Astronomy Data Centre (CADC/NRC/CSA). Based on observations collected at the European Southern Observatory under ESO programme 094.B-0771 obtained from the ESO Science Archive Facility with DOI https://doi.eso.org/10.18727/archive/42.
@@
The HST imaging data is publicly available at the Hubble Legacy Archive (https://hla.stsci.edu/) under the programs 11602 (PI: Sahar Allam) and 12266 (PI: Anna Quider). The MUSE data is available at ESO Science Archive Facility (http://archive.eso.org/scienceportal/) under the program-ID 094.B-0771 (PI: Bethan James). Data relative to Fig. 10 are available at https://github.com/remco-space/Black-Hole-Mass-compilation. Additional data generated in this research will be available upon request to the corresponding author.

### v3

@@ 4.2.1 Fiducial model perturbations
There are three other sources of systematics linked to the fiducial model. First, the alignment of the velocity ellipsoid, which is assumed to align with the spherical coordinate system. Second, the source plane pixelisation grid, which is based on a Voronoi tessellation. And third, the use of the prior on the total projected mass within the Cosmic Horseshoe Einstein ring.
In model M6, we test the effect of a cylindrical alignment (Cappellari, 2008) for the velocity ellipsoid, while in model M7, we change the source plane pixelisation from a Voronoi to a Delaunay tessellation. In both cases, we recover SMBH masses that agree with the fiducial model within the $1\sigma$ level.
Finally, we assess the impact of the prior on the total projected mass within the Einstein radius of s2 by removing this constraint in model M8. This model leads to $\log_{10}(M_{\text{BH}}/M_{\sun})=10.51^{+0.07}_{-0.09}$, showing that the prior has a minimum impact on the determination of the SMBH mass.
@@ 4.2.2 More flexible mass models
We now explore more flexible mass models to assess whether increased freedom in the mass distribution can account for the high SMBH mass inferred in our fiducial model.
@@
Further, model M10 incorporates additional freedom by assigning each luminous Gaussian its own anisotropy parameter. The DM halo is modelled as a spherical NFW profile with the scale radius fixed to the fiducial model. With this configuration, we recover an SMBH mass of $\log_{10}(M_{\text{BH}}/M_{\sun})=10.55^{+0.10}_{-0.07}$.
@@
  • Model M11 assuming the main mass-concentration relation, yields an SMBH mass of $\log_{10}(M_{\text{BH}}/M_{\sun})=10.15^{+0.17}_{-0.30}$.
@@
These results, once more, highlight the robustness of the fiducial model, as they agree with the fiducial model within the uncertainties.
4.3 Is an SMBH necessary?
So far, we only fit models with a presence of an SMBH, and despite the many variations of the mass profile, our results are fairly consistent between each other. However, one might question whether an SMBH is necessary to explain the observed data. To answer this question, we fitted model M14 using the same fiducial mass model, but without including the SMBH component.
Fig. 9 shows the highest-likelihood lens model in the upper panels, and the dynamical model in the bottom panels for model M14. Qualitatively, the fiducial lens model and the M14 lens model are very similar. Both are able to reproduce the observed data with similar residuals, and to reconstruct the source with close morphologies. On the other hand, the dynamical model fit the data poorly, especially at the central regions, where the SMBH presence is expected to be more relevant.
Figure 9: Model results of mass configuration M14, which differs from the fiducial model by not including an SMBH. The upper panels display, left to right, the highest-likelihood lens model, the normalised residuals, and the source reconstruction. The bottom panels are, left to right, the median dynamical model, the normalised residuals, and the radial kinematic profile with the median model. All top images are in units of electrons per second.
We can also use the Bayesian evidence of the models to quantitatively assess the need of an SMBH. The third column of Table 3 summarises the relative Bayesian evidence ($\Delta\ln{\mathcal{Z}}$) for all models considered in this work, calculated with respect to the fiducial model. Comparing the fiducial model M1 and model M14, which excludes the SMBH, we find a difference in the Bayesian evidence of $\Delta\ln{\mathcal{Z}}=16.31$. This corresponds to a statistical preference exceeding $5\sigma$ in favour of the SMBH [^13] This assumes equal prior model probabilities..
We also attempt to fit model M4 without the inclusion of an SMBH (M15) to evaluate whether the central mass-to-light component could replicate the SMBH’s contribution. In this configuration, the central Gaussian component is assigned its own $\Upsilon_{\star}$, potentially steepening the stellar mass density profile near the galaxy’s centre and compensating for the absence of the SMBH. However, as with model M14, this approach failed to reproduce the observed kinematical data accurately. A comparison of the Bayesian evidence for this model against the fiducial model yields $\Delta\ln{\mathcal{Z}}=11.17$, strongly favouring the fiducial model. This difference corresponds to a $5\sigma$ detection of the SMBH.
@@ 4.4 The SMBH mass
In Table 3, we present the SMBH masses and relative (in respect to the fiducial model) Bayesian evidences for all the mass models in this work. The highest Bayesian evidence is associated with model M3, which incorporates a variable anisotropy profile. A comparison between the fiducial model M1 and model M3 yields $\Delta\ln{\mathcal{Z}}=-3.48$, corresponding to a $0.9\sigma$ preference to M3. This evidence difference is not decisive, and we do not have sufficient prior knowledge to say if one model should be astrophysically preferred over the other. Additionally, measuring the anisotropy profile is notoriously challenging due to its degeneracies with other parameters and sensitivity to data quality, which is why we adopted the fiducial model for its simplicity and robustness.
Given this result, we adopt the SMBH mass inferred by the fiducial model as our final value, and we take the scatter between the alternative mass as an estimate of the systematic uncertainty. Using the standard deviation across all SMBH mass measurements, our final inference is $\log_{10}(M_{\text{BH}}/M_{\sun})=10.56^{+0.07}_{-0.08}\pm(0.12)^{\text{sys}}$ at $1\sigma$ level, confirming the detection of an UMBH in the Cosmic Horseshoe main-lens galaxy.
@@ 4.5 The role of the radial image
To explore this, we performed dynamical-only modelling for all the mass models listed in Table 3, and we show the resulting SMBH mass measurements in Table 4.
Table 4: Dynamical models only SMBH results. Values are the median and $1\sigma$ uncertainties. Models M7 and M8 are not applicable.
@@
Comparing the jointly results with those obtained through dynamical-only modelling, we find that the latter, in general, yields more massive SMBH estimates. Additionally, the error bars for the dynamical-only models are larger, which is expected due to the smaller number of data points in the kinematic map. Using the same criteria as before to determine the final SMBH mass, the dynamical-only SMBH mass is $\log_{10}(M_{\text{BH}}/M_{\sun})=10.72^{+0.10}_{-0.13}\pm(0.30)^{\text{sys}}$ at $1\sigma$ level.
These findings highlight the important role of lensing information in constraining the SMBH mass, particularly by limiting how massive the SMBH can be. This is especially significant in the context of direct SMBH mass determinations in intermediate and high-redshift systems, where IFU data often suffers from suboptimal spatial resolution and SNR. When the radial image is well-resolved and has sufficient SNR, the lensing effect is sufficient to effectively constrain $M_{\text{BH}}$, as demonstrated by Nightingale et al. (2023). On the other hand, when image quality is less favourable — reflecting the challenges of observing more distant systems — the integration of dynamical and lensing data becomes essential for reliable SMBH mass measurements.
@@ 5.1 SMBHs and the $M_{\text{BH}}-\sigma_{e}$ relation
In Fig. 10, we put the SMBH of the main deflector in the Cosmic Horseshoe lens system in the context of the $M_{\text{BH}}-\sigma_{e}$ relation from van den Bosch (2016). The SMBH reported here is among the most massive SMBHs ever detected, so is the galaxy that hosts it: the measured effective velocity dispersion of $\sigma_{e}=366\pm 6$ km s-1. Other SMBHs exceeding $10^{10}M_{\sun}$ — and thus also classified as UMBHs — with comparable $\sigma_{e}$ have been found in Holm 15A (Mehrgan et al., 2019) and NGC 4889 (McConnell et al., 2011), both of which are nearby BCGs. The lensing galaxy of the Cosmic Horseshoe system is unique in that it lies at $z_{l}=0.44$ and lacks comparably massive companion galaxies — it is possibly the central member of a fossil group or loose cluster (Ponman et al., 1994; Dye et al., 2008).
@@
One possible scenario involves binary SMBH scouring, a process that can occur during the merger of massive galaxies and is more likely in the central galaxy of a group or cluster. In this process, the binary SMBHs dynamically expel stars from the central regions of the merged galaxy, effectively reducing the stellar velocity dispersion while leaving the SMBH mass largely unchanged (e.g., Thomas et al., 2014, 2016; Dullo, 2019). Another possible scenario involves AGN feedback processes, where powerful outflows and jets may quench star formation and alter the galaxy’s central structure, decoupling the growth of the SMBH from the host galaxy’s stellar kinematics. Strong AGN feedback can also transfer energy to the DM and stellar components, modifying the central surface brightness profile and mimicking the presence of a core (see discussion in Mehrgan et al., 2019). A third scenario posits that such UMBHs could be remnants of extremely luminous quasars, which experienced rapid SMBH accretion episodes in the early Universe (McConnell et al., 2011; Wu et al., 2015).
@@
5.2 Other astrophysical implications
Beyond the determination of the SMBH mass, we can infer other physical properties of the main deflector.
@@ 5.2 Other astrophysical implications
The inner DM density slope is another noteworthy quantity, as it provides critical insights into the interaction between baryons and DM (e.g., Gnedin et al., 2004; Petit et al., 2023). Early N-body DM-only simulations suggested that haloes are well described by the NFW profile (Navarro et al., 1997). However, the inclusion of baryonic components, especially feedback processes (e.g., Di Cintio et al., 2014; Jackson et al., 2024), has been shown to alter the DM distribution within galaxies. These modifications may be linked to longstanding issues such as the “cusp-core” problem (see Del Popolo & Le Delliou, 2021, for a review).
@@
Fig. 11 compares the surface mass density profiles along the semi-major axis for these three models. All models exhibit strong agreement, particularly in the inner regions. The most notable deviation occurs in the outermost region of the stellar density profile for model M9, but this remains within the $1\sigma$ uncertainty. The larger uncertainties in model M9 can be explained by the greater freedom in its mass profile, which allows for simultaneous variations in the stellar mass-to-light ratio, stellar anisotropy, and DM inner slope. Additionally, the galaxy is already DM dominated before reaching the effective radius.
Figure 11: Surface mass density profile for three models — fiducial (upper) and other two that uses an gNFW profile: M2 (middle) and M9 (lower). The M2 model resembles the fiducial model but with the DM inner slope as a free parameter, while M9 incorporates both a variable mass-to-light ratio and a variable anisotropy profile. Blue lines represent the stellar component, purple lines represents the DM component, and in black we show the total (DM+stellar+BH) surface density. The shaded regions indicate the $1\sigma$ credible intervals for each component. The horizontal dashed lines represents the effective radius and the Einstein radius.
We also employed a gNFW halo in models M11-M13, but incorporating a mass-concentration relation to determine the scale radius. The recovered inner DM slopes for these models were $\gamma_{\text{DM}}=1.29^{+0.05}_{-0.04}$, $\gamma_{\text{DM}}=1.34^{+0.07}_{-0.07}$, and $\gamma_{\text{DM}}=1.12^{+0.03}_{-0.04}$ respectively. These slopes are steeper than an NFW-like profile. However, it is important to note that these steeper slopes do not result in better fits or higher Bayesian evidences. On the contrary, models M11-M13 exhibit lower evidences (see Table 3) compared to the other gNFW models that recovered an NFW-like slope. Additionally, when comparing these models with the fiducial model, the fiducial model remains slightly preferred.
This topic was also investigated by Schuldt et al. (2019) in their analysis of the Cosmic Horseshoe system. They reached a similar conclusion, finding that allowing more flexibility in the inner density slope of the DM halo did not lead to a significant improvement in the fit. Their results, like ours, suggest that the halo is either NFW-like or deviates only slightly from this profile, indicating that baryonic processes either do not significantly alter the inner DM halo structure in this galaxy, or cancel each other out.
@@ 6 Conclusions
In this work, we investigated the inner structure of the Cosmic Horseshoe lensing galaxy by applying a self-consistent model to both the radial arc and the stellar dynamics of the main deflector, using the mass of the main ring as a prior. Our fiducial model constrained the mass of the SMBH at the centre of the main-lens to be $\log_{10}(M_{\text{BH}}/M_{\sun})=10.56^{+0.07}_{-0.08}\pm(0.12)^{\text{sys}}$, thus an UMBH. We rigorously tested a variety of systematics, including uncertainties in the mass profile and modeling choices, but all models consistently converged to the fiducial value. A Bayesian model comparison revealed a $5\sigma$ detection of the SMBH relative to a model without a its contribution, reinforcing our results.
@@
Such analysis present here was only possible due to the observation of a radial arc. Radial arcs, like the one studied here, are expected to become increasingly common. The Euclid mission is expected to discover hundreds of thousands of lenses over the next five years (Collett, 2015; Euclid Collaboration et al., 2025), while the Extremely Large Telescope (ELT) will revolutionize our ability to conduct detailed dynamical studies. The combination of lensing and dynamics will soon provide an unprecedented sample of galaxies, offering exciting insights into stellar populations, DM halos, and SMBHs across a broader redshift range than ever before. This new era of discovery promises to deepen our understanding of galaxy evolution and the interplay between baryonic and DM components.
@@ Acknowledgements
We thank the anonymous referee for the detailed feedback. This work has been supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq). This research was supported by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) through grant 88887.936450/2024-00. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536). TEC is funded by the Royal Society through a University Research Fellowship. CF ackwoledges funding from CNPq and the Rio Grande do Sul Research Foundation (FAPERGS) through grants CNPq-315421/2023-1 and FAPERGS-21/2551-0002025-3. ACS acknowledges funding from CNPq and FAPERGS through grants CNPq-314301/2021-6 and FAPERGS/CAPES 19/2551-0000696-9. The authors acknowledge the National Laboratory for Scientific Computing (LNCC/MCTI, Brazil) for providing HPC resources of the SDumont supercomputer, which have contributed to the research results reported within this paper. URL: http://sdumont.lncc.br. This work made use of the CHE cluster, managed and funded by COSMO/CBPF/MCTI, with financial support from FINEP and FAPERJ, and operating at the Javier Magnin Computing Center/CBPF. Based on observations made with the NASA/ESA Hubble Space Telescope, and obtained from the Hubble Legacy Archive, which is a collaboration between the Space Telescope Science Institute (STScI/NASA), the Space Telescope European Coordinating Facility (ST-ECF/ESA) and the Canadian Astronomy Data Centre (CADC/NRC/CSA). Based on observations collected at the European Southern Observatory under ESO programme 094.B-0771 obtained from the ESO Science Archive Facility with DOI https://doi.eso.org/10.18727/archive/42.
@@
The HST imaging data are publicly available at the Hubble Legacy Archive (https://hla.stsci.edu/) under the programs 11602 (PI: Sahar Allam) and 12266 (PI: Anna Quider). The MUSE data are available at ESO Science Archive Facility (http://archive.eso.org/scienceportal/) under the program-ID 094.B-0771 (PI: Bethan James). Data relative to Fig. 10 are available at https://github.com/remco-space/Black-Hole-Mass-compilation. Additional data generated in this research will be available upon request to Carlos Melo: carlos.melo@ufrgs.br.